OpenStudy (anonymous):

A theater company has raised $810.75 by selling 15 floor seat tickets. Each ticket costs the same. Part A: Write an equation with a variable that can be solved to correctly find the price of each ticket. Explain how you created this equation. Part B: Solve your equation in Part A to find the price of each floor seat ticket. How do you know your solution is correct?

OpenStudy (automaticloveletter):

$810.75 is the total price that they raised 15 floor seat tickets were sold 15x = 810.75 How to solve? inverse operation so divide 15/15 (cancels out) 810.75/15=54.05 Thus, each floor seat ticket costs $54.05 Hope I helped! :)

OpenStudy (anonymous):

which one is part a and which one is part b

OpenStudy (automaticloveletter):

Part A: Write an equation with a variable that can be solved to correctly find the price of each ticket. Explain how you created this equation. $810.75 is the total price that they raised 15 floor seat tickets were sold 15x = 810.75 Part B: Solve your equation in Part A to find the price of each floor seat ticket. How do you know your solution is correct? How to solve? Using inverse operations to divide 15/15 (cancels out) 810.75/15=54.05 Thus, each floor seat ticket costs $54.05 Check: 54.05*15=810.75

OpenStudy (anonymous):

which one is part a and which one is part b

OpenStudy (anonymous):

Thanks

OpenStudy (automaticloveletter):

You're welcome :)

OpenStudy (tabdo1226):

Go real madrid boo barcelona

OpenStudy (texaschic101):

15x = 810.75 is part A....with x representing the cost of each ticket so basically this breaks down to 15 tickets at $x = 810.75 part B : solve it 15x = 810.75 -- divide both sides by 15 x = 54.05 and you can check your answer by subbing it back into the original equation to see if it is true 15x = 810.75 15(54.05) = 810.75 810.75 = 810.75 (correct).....so that proves your answer